IVE HOMEWORK Saved Help Lenow Drug Stores and Hall Pharmaceuticals are competitors in the discount drug chain store business. The separate capita structures for Lenow and Hall are presented here. Hall Debt @ 10 Common stock, $10 par Total Common shares $100,000 Debt @ 10% 200,000 Common stock, $10 par 300,000 Total $200,000 100,000 300,000 10,000 20,000 Common shares a. Complete the following table given earnings before interest and taxes of $20,000, $30,000, and $120,000. Assume the tax rate is 30 percent. (Round your answers to 2 decimal places.) What is the relationship between the EPS of the two firms? EBIT Total Assets | S 20,000$300,000 $ 30,000$300,000 120,000 $ 300,000 EBIT/TA | % | Lenow EPS Hall EPS
b-1. What is the EBIT/TA rate when the firms have equal EPS? EBIT/TA rate b-2. What is the cost of debt? Cost of debt b-3. State the relationship between earnings per share and the level of EBIT the cost of debt. EPS is unaffected by financial leverage when the pre-tax return on assets (EBIT/TA)
b-3. State the relationship between earnings per share and the level of EBIT EPS is unaffected by financial leverage when the pre-tax return on assets (EBIT/TA) the cost of debt c.If the cost of debt went up to 12 percent and all other factors remained equal, what would be the break-even level for EBIT? Break-even level
Answer #1
a)
EBIT TA EBIT/TA LENOW EPS HALL EPS RELATIONSHIP EPS CALCULATION:
20000 300000 0.07 0.5 0 as the EBIT rises, EBIT - INTEREST / NUMBER OF SHARES
30000 300000 0.1 1 1 the EPS of Hall increases at
120000 300000 0.4 5.5 10 higher rate
b1) When the firm's have equal EPS, the EBIT/TA is 10%.
b2) The cost of debt is 10% for both the firms
b3) sorry, have not got the statement
c) Break even level for EBIT when cost of debt went to 12%.
LENOW HALL
EBIT - 12000 / 20000 = EBIT - 24000 /10000
EBIT - 12000 = (EBIT - 24000)*2
BEL of EBIT = 48000-12000 = $36000